OpenStudy (anonymous):

Factor trinomial a^2+9a+20 I got (a+5)(a+4) Check please

OpenStudy (anonymous):

yeah that's right multiply it back out to check: (a+5)(a+4)=a^2+4a+5a+20=a^2+9a+20

OpenStudy (amistre64):

5(4) = 20 5+4 = 9 ... youre good :)